# Eleanor Wiske Dillon Sonia Jaffe Microsoft Research Microsoft Research Nicole Immorlica Christopher T. Stanton Microsoft Research Harvard Business School, NBER, and CEPR

### v1

@@ Appendix A Experiment Details
More details on the study design and implementation are available in Dillon et al. (2025). At the time the experiment was run, firms could purchase at most 300 M365 Copilot licenses. The Microsoft marketing team recruited firms to join the experiment with the understanding that at least 50 of their licenses would be allocated at random between workers. Firms were asked to recruit at least twice the number of workers for the study as the number of Copilot licenses that would be randomly assigned. For example, a firm that chose to allocate 50 of their 300 licenses at random would need to recruit at least 100 workers to join the experiment, while a firm that chose to allocate 150 licenses at random would need to recruit 300 workers to join. Some firms had more than 400 workers participate in the study, while a few ended up with slightly less than 100. Each individual firm selected which workers they wanted to participate; some focused on a single department or work area while others pulled workers from many parts of the company. Firms were responsible for all communications with the participating workers, including obtaining necessary consent. Some firms provided additional information about the recruited workers such as department or job title. While the sample of workers with this additional information is too small for meaningful subgroup analyses, it confirms that the participating workers are generally early to mid-career knowledge workers such as analysts, project managers, or professional service providers.
Our analysis sample has 56 firms across a range of industries, as shown in Figure 1(a). There were 4 additional firms that agreed to participate in the study but did not follow through with randomization; these firms are excluded from the analysis sample. The experiment includes workers from many countries, with the majority based in the United States or Europe.
@@
For all firms, our data sample begins on May 5, 2023, providing at least 4 months of data before any Copilot licenses were allocated. Firms varied in when they allocated licenses, with the first starting in late September 2023, and the latest in April 2024. Figure 1(b) illustrates the number of firms actively maintaining randomization over the study window. We collected data on the workers at each firm for a maximum of six months after licenses were allocated. Six months is also the modal length of the post period across firms. A small number of firms have shorter treatment periods because they explicitly exited the experiment prior to the six-month post-period. In all analyses, we drop outcomes for the weeks ending December 29, 2023 and January 5, 2024 as work activity was low across all workers during this holiday period.
@@ Appendix B Data Details
We analyzed millions of Microsoft 365 usage signals (telemetry), for both the treated and designated control workers, from product telemetry data that Microsoft collects. To maintain workers’ and firms’ privacy, we do not observe anything about the content produced by these workers, so our focus is on how workers allocate their time and interact with each other rather than on the quality of their work.
For some analyses, we restrict our estimation to a subsample of workers who used a particular application in question. While virtually every worker in our sample used Outlook regularly, fewer workers used Word, indicating that many workers likely do not have jobs that require document creation in Word. For results that examine document creation or interactions, we subset the analysis to workers with higher Word use in the pre-period, indicated by above-median time spent in Word (.27 hours/week) and number of documents originated as the ‘primary editor’ (.19 documents/week). For Teams meetings, we drop five firms that were not regular users of Teams prior to the experiment because they relied on a different video meeting program.
@@ B.1 Outcome Metrics
We construct weekly metrics capturing each worker’s email, document editing, and meeting behaviors. Metrics are winsorized to the 95th percentile, except Word document count metrics, which are winsorized to the 99th percentile due to fewer weekly observations.
@@
    For email, we aggregate the signals of various ‘reading actions’ individuals take into ‘Outlook sessions’: blocks of time where they don’t go more than 15 minutes without reading an email.
@@
    • Outlook out-of-hours total session time The subset of Outlook session time that occurred outside the individual’s “working day” as defined as the 8 hours in which they were most active during one month of pre-period.
@@
    • Unique email threads replied to For emails received that week that are replied to within 7 days, the number of unique email threads/conversations.
    • Time to reply from email delivery For emails received that week that are replied to within 7 days, the average length of time it takes to reply to an email.
@@
    • Total emails in email threads replied to For emails received that week that are replied to within 7 days, the total number of emails in the threads/conversations.
@@
    • Teams meeting total session time The time spent in scheduled Teams meetings (excluding spontaneous calls) in a week, measured in hours.
    • Share of meetings attended 5 min late/early The share of scheduled Teams meetings joined either $\geq$5 minutes after the Teams meeting was started or left $\geq$5 minutes before the meeting was ended.
    • Recurring Teams meetings total session time The time spent in Teams meetings marking as recurring meetings, measured in hours.
@@
    • Across metrics, a worker is a primary editor if they make the most edits on a document; a document is completed when 90% of the total edits to the document have been completed. [^14] Some documents show stray edits long after most active work is done, which may be accidental. This 90% cutoff aligns most often with other indicators of completion such as attaching to an email or printing to pdf.
    • Word total session time The time spent in Word sessions, identified as blocks of actions with $<15$ minutes gap, measured in hours.
@@
    • Avg. collaborative document time to complete The same as the average time to complete, but only for documents with multiple collaborators.
@@
    • Others’ documents read/edited before completed The number of documents read or edited in a week for which the target worker is not the primary editor.
@@ B.2 Coworkers
We also use the telemetry data to identify the closest coworkers of each person in the experiment, based on a weighted sum of their joint meetings, calls, emails, chats, and document collaborations. For each metric, we divide by the firm-specific average. We add these normalized numbers together for an interaction score and call someone a “coworker” if the score is greater than 8.0 (If they add the average amount of interaction, the score would be 5.0 since there are five metrics.) If an individual has more than 15 coworkers, we take only the 15 with the highest interaction score. The average worker has 9 coworkers. We can then calculate what share of an individual’s coworkers have a Copilot license. Figure A2 shows the distribution of that share.
@@ C.1 Adoption
Figure A3 shows some additional Copilot usage patterns. Figure 3(a) shows the variation across firms in usage over time. Figure A3 shows a binscatter of the mean weekly Copilot actions taken in each app among users using some form of Copilot that week by the firm-level share of users that are active in an average week. Similar to Figure 1(b), we see that the usage intensity grows with firm usage much more for Teams than for Outlook (with Word in between).
@@
Emails read -0.0003∗∗ -0.0003∗∗ -0.0002∗∗ -0.0004∗∗
($7.91\times 10^{-5}$) ($7.95\times 10^{-5}$) ($8.33\times 10^{-5}$) ($8.09\times 10^{-5}$)
Unique email threads replied to -0.0014 -0.0014 -0.0011 -0.0010
@@
Teams meeting total session time 0.0109∗∗ 0.0108∗∗ 0.0107∗∗ 0.0127∗∗
@@
Teams meetings attended 0.0068∗∗ 0.0067∗∗ 0.0040 0.0050∗∗
@@
Word total session time 0.0099 0.0107 0.0132 0.0213∗∗
@@
Others’ documents read/edited before completed 0.0590∗∗ 0.0574∗∗ 0.0340∗∗ 0.0570∗∗
@@
Number of email recipients -0.0009 -0.0009 -0.0008 -0.0009
@@
Number of unique people met 1-1 (total) -0.0004 -0.0004 0.0003 -0.0004
@@
Share of coworkers with Copilot license 0.0439 0.0528 0.0988∗∗ 0.0384
@@
Fixed-effects
Firm Yes Yes Yes
Industry Yes Yes
Fit statistics
Observations 2,718 2,718 2,718 3,167 3,167 3,167 2,718
R2 0.09294 0.09412 0.23525 0.18860 0.18388 0.05045 0.13906
Within R2 0.05276 0.00578 0.09072
Significance: ** $p<0.01$, * $p<0.05$
Note: This table shows OLS regressions results using individuals’ pre-experiment behaviors, share of coworkers with a Copilot license, and firm and industry fixed effects to predict mean post-period Copilot weekly usage among treated workers. All variables are averages over the pre-period other than “number of unique people met 1-1,” which is a total over a fixed pre-period.
Table A2: Predicting Weekly Copilot Use
Emails read -0.0003∗∗ -0.0003∗∗ -0.0002∗∗ -0.0004∗∗
($5.5\times 10^{-5}$) ($5.53\times 10^{-5}$) ($5.75\times 10^{-5}$) ($5.61\times 10^{-5}$)
Unique email threads replied to -0.0013∗∗ -0.0014∗∗ -0.0010 -0.0009
@@
Teams meeting total session time 0.0097∗∗ 0.0096∗∗ 0.0105∗∗ 0.0121∗∗
@@
Teams meetings attended 0.0074∗∗ 0.0074∗∗ 0.0041∗∗ 0.0053∗∗
@@
Word total session time 0.0127∗∗ 0.0131∗∗ 0.0134 0.0218∗∗
@@
Others’ documents read/edited before completed 0.0498∗∗ 0.0491∗∗ 0.0341∗∗ 0.0534∗∗
@@
Number of email recipients -0.0007∗∗ -0.0007∗∗ -0.0008∗∗ -0.0008∗∗
@@
Number of unique people met 1-1 (total) -0.0002 -0.0002 0.0002 -0.0003
(0.0003) (0.0003) (0.0003) (0.0003)
Share of coworkers with Copilot license 0.0236 0.0546∗∗ 0.1008∗∗ 0.0361
(0.0169) (0.0172) (0.0164) (0.0169)
@@
Firm Yes Yes Yes
Industry Yes Yes
@@ C.2 Outcomes
Note: These plots show the coefficients from regressions similar to Equations (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates, where the $\beta$’s are allowed to vary by firm’s month from Copilot adoption. Vertical lines show 95% confidence intervals based on Newey-West standard errors with a lag of 9 weeks. Horizontal dashed lines give the average estimated effect from Table 2.
Table A3 shows the same analysis as Table 2 for some additional outcome variables. The decrease in emails read individually is unsurprising if Copilot is summarizing email threads (or all new emails) for people. Total emails replied to is another measure of reply activity where while the difference is statistically significiant, it is practically very small ($<2\%$).
We also look at a bunch of different types of meetings. While we do see some increases for big meetings, the event studies in Figure LABEL:fig:app_event_studies2 show that these increases only happened at the very start of the study period and seem likely due to the Copilot trainings that many companies did.
We also ran a balance test of workers’ pre-period behavior between treatment and control groups and found no significant differences across all outcomes.
Table A3: Effect of Copilot
@@
Total emails in email threads replied to 3.79 -0.036** -0.073** 6,111 285,332
@@
Total Teams meetings attended 9.92 0.11 0.23 5,698 299,922
@@
Total recurring meetings 4.14 0.046 0.091 5,698 254,319
@@
Small short meetings 6.81 0.038 0.074 5,698 254,319
@@
Small long meetings 1.15 -0.004 -0.007 5,698 254,319
(0.80) (0.010) (0.019)
[…]
Others’ documents read/edited 0.44 0.030 0.053 2,150 91,213
@@
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Column 1), the intent-to-treat coefficients for the effect of Copilot (Column 2) and the LATE coefficients (Column 3). Cross-user standard deviations are estimates of the standard deviation of the random intercept in the model $Y_{it}=\mu_{i}+\epsilon_{it}$.
@@
Note: These plots show the coefficients from regressions similar to Equations (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates, where the $\beta$’s are allowed to vary by firm’s month from Copilot adoption. Vertical lines show 95% confidence intervals based on Newey-West standard errors with a lag of 9 weeks. Horizontal dashed lines give the average estimated effect from Table A3.
Table A4: Pre-period Balance Test
Metric Control Mean Treated Mean $q$-value
Outlook total session time 11.59 11.48 1
Total concentration time 24.55 24.67 1
Outlook out-of-hours session time 2.11 2.05 1
Unique email threads replied to 13.07 13.36 1
Time to reply from email delivery 14.63 14.89 1
Teams meeting total time 5.16 5.26 1
% Teams meetings attended late/early 0.29 0.30 1
Recurring Teams meetings time 2.28 2.35 1
Word total session time 1.37 1.33 1
Avg. document TTC 181.45 170.02 1
Avg. collaborative document TTC 270.14 286.12 1
Avg. non-collaborative document TTC 149.45 138.37 1
Completed documents 0.79 0.76 1
Emails read 156.26 154.92 1
Total emails in email threads replied to 3.79 3.79 1
Total Teams meetings attended 9.83 10.00 1
Total recurring meetings 4.07 4.21 1
Big short meetings 1.03 1.06 1
Big long meetings 0.86 0.88 1
Small short meetings 6.72 6.89 1
Small long meetings 1.14 1.17 1
Others’ documents read/edited 0.44 0.43 1
Number of email recipients 4.08 4.59 1
Number of unique people met 1-1 (total) 15.25 15.53 1
C.3 Splitting Sample by Firm Adoption
Table A5 is similar to Table 3 in the paper, but instead of splitting individuals by the share of their coworkers who had a Copilot licenses, it splits by whether the average Copilot usage rate at the firm was above or below the median of 45%. As one would expect the OLS coefficients are smaller at the firms where people used Copilot less – less usage leads to a smaller effect.
The IV results in Table A5 are somewhat hard to interpret because they are basically rescaling the ITT (OLS) results by the share of workers using Copilot while simultaneously splitting on that share. As shown in Figure 1(b), at low-usage firms a larger share of those using Copilot are using Outlook Copilot, which probably explains the larger IV coefficients for Outlook variables. It is not that Copilot usage actually had a larger effect at the lower-usage firms, but we are in some sense over-measuring the relevant (i.e. Outlook) Copilot usage at the high-usage firms.
Table A5: Effect of Copilot - above and below median firm usage of Copilot
Metric Pre-period OLS (SE) IV (SE) $\mathbf{n}$
mean (SD)
Above Below Above Below Above Below Above Below
Outlook total session time 11.32 (6.21) 11.80 (6.60) -1.47** (0.096) -1.21** (0.099) -2.46** (0.16) -4.22** (0.35) 156,648 124,635
Total concentration time 24.72 (5.97) 24.47 (6.45) 0.94** (0.083) 0.86** (0.089) 1.32** (0.14) 3.19** (0.33) 137,014 110,893
Outlook out-of-hours session time 2.00 (1.59) 2.19 (1.71) -0.34** (0.026) -0.25** (0.027) -0.56** (0.043) -0.88** (0.094) 156,648 124,635
Unique email threads replied to 13.12 (10.87) 13.34 (12.11) -0.31* (0.11) -0.13 (0.12) -0.50* (0.18) -0.41 (0.37) 181,898 143,574
Time to reply from email delivery 14.56 (4.50) 15.03 (5.18) -0.38** (0.12) -0.21 (0.14) -0.61** (0.19) -0.63 (0.43) 160,590 124,742
[…]
Teams meeting time 5.56 (3.23) 4.73 (3.29) 0.030 (0.045) 0.056 (0.050) 0.048 (0.072) 0.17 (0.15) 122,568 86,186
% Teams meetings attended late/early 0.29 (0.097) 0.30 (0.12) 0.010** (0.002) 0.002 (0.003) 0.015** (0.004) 0.005 (0.009) 107,111 71,107
Recurring Teams meetings time 2.53 (1.72) 2.02 (1.60) 0.018 (0.021) 0.006 (0.023) 0.028 (0.033) 0.019 (0.068) 148,836 105,483
[…]
Word total session time 1.42 (0.96) 1.24 (0.87) 0.044 (0.030) -0.025 (0.034) 0.066 (0.045) -0.068 (0.093) 71,076 43,787
Avg. document TTC 176.35 (84.33) 174.26 (80.63) -9.05 (6.67) -3.13 (9.04) -12.85 (9.46) -7.72 (22.25) 22,877 13,646
Avg. collaborative document TTC 282.10 (115.79) 271.15 (111.07) -37.51 (16.59) -30.61 (26.14) -52.44 (23.14) -65.23 (55.66) 7,429 3,604
Avg. non-collaborative document TTC 141.39 (80.15) 147.65 (74.08) -5.77 (6.51) -0.51 (8.63) -8.18 (9.23) -1.27 (21.33) 19,111 11,864
Completed documents 0.78 (0.73) 0.77 (0.77) 0.047* (0.021) -0.003 (0.028) 0.070* (0.032) -0.008 (0.076) 56,358 34,855
Completed collaborative documents 0.16 (0.25) 0.12 (0.18) 0.007 (0.007) 0.017 (0.008) 0.011 (0.011) 0.045 (0.023) 56,358 34,855
Significance: ** $q<0.01$, * $q<0.05$
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Columns 1 & 2), the intent-to-treat coefficients for the effect of Copilot (Columns 3 & 4), and the LATE estimates (Columns 5 & 6) separately for workers at firms who had above or below median Copilot usage in the post-period. All time metrics in hours. A worker is a primary editor if they make the most edits on a document. Time to complete (TTC) and completed document counts are calculated only for documents where the target worker is the primary editor.

### v2

@@ Appendix A Experiment Details
More details on the study design and implementation are available in Dillon et al. (2025). For example, a firm that chose to allocate 50 of their 300 licenses at random would need to recruit at least 100 workers to join the experiment, while a firm that chose to allocate 150 licenses at random would need to recruit 300 workers to join. Some firms had more than 500 workers participate in the study, while one firm had only 68 workers participate. Each individual firm selected which workers they wanted to participate; some focused on a single department or work area while others pulled workers from many parts of the company. Firms were responsible for all communications with the participating workers, including obtaining necessary consent. Some firms provided additional information about the recruited workers such as department or job title. While the sample of workers with this additional information is too small for meaningful subgroup analyses, it confirms that the participating workers are generally early to mid-career knowledge workers such as analysts, project managers, or professional service providers.
Our analysis sample has 66 firms across a range of industries, as shown in Figure 1(a). There were 2 additional firms that agreed to participate in the study but did not follow through with randomization; these firms are excluded from the analysis sample. The experiment includes workers from many countries, with the majority based in the United States or Europe.
@@
For all firms, our data sample begins on May 5, 2023, providing at least 4 months of data before any Copilot licenses were allocated. Firms varied in when they allocated licenses, with the first starting in September 2023, and the latest in April 2024. Figure 1(b) illustrates the number of firms actively maintaining randomization over the study window. We collected data on the workers at each firm for a maximum of 24 weeks after licenses were allocated. 24 weeks is also the modal length of the post period across firms. 5 firms have shorter treatment periods because they explicitly exited the experiment prior to the 24-week post-period. In all analyses, we drop outcomes for the weeks ending December 29, 2023 and January 5, 2024 as work activity was low across all workers during this holiday period.
Table A1 shows the balance tests for pre-experiment differences between the treatment and control groups. There are no significant differences across all outcomes.
Table A1: Pre-period Balance Test
Main Outcomes Control Mean Treated Mean $p$-value
Outlook total session time 11.73 11.58 0.45
Total concentration time 26.10 26.27 0.37
Outlook out-of-hours session time 2.22 2.17 0.42
Unique email threads replied to 13.52 13.59 0.84
Time to reply from email delivery 16.27 16.58 0.05
Teams meeting total time 5.17 5.28 0.23
% Teams meetings attended late/early 0.30 0.31 0.13
Recurring Teams meetings time 2.28 2.35 0.15
Word total session time 1.67 1.58 0.17
Avg. document TTC 191.11 182.14 0.26
Avg. collaborative document TTC 276.04 299.36 0.10
Avg. non-collaborative document TTC 163.61 152.75 0.14
Completed documents 0.78 0.74 0.30
Completed collaborative documents 0.14 0.13 0.28
Additional Metrics Control Mean Treated Mean $p$-value
Emails read 158.03 157.00 0.78
Total emails in email threads replied to 3.83 3.83 0.77
Total Teams meetings attended 9.84 10.01 0.40
Total recurring meetings 4.06 4.19 0.15
Big short meetings 1.08 1.09 0.87
Big long meetings 0.89 0.90 0.71
Small short meetings 6.63 6.83 0.14
Small long meetings 1.19 1.23 0.17
Others’ documents read/edited 0.41 0.39 0.34
Number of email recipients 4.09 4.49 0.05
Number of unique people met 1-1 (total) 13.90 14.15 0.39
@@ Appendix B Data Details
We analyzed hundreds of millions of Microsoft 365 usage signals (telemetry), for both the treated and designated control workers, from product telemetry data that Microsoft collects. To maintain workers’ and firms’ privacy, we do not observe anything about the content produced by these workers, so our focus is on how workers allocate their time and interact with each other rather than on the quality of their work.
One limitation of this study is that we only observe activity within the Microsoft Office products we can track. We are, for example, measuring time spent in meetings through Teams, not time spent in meetings overall. For email outcomes, this distinction appears unimportant. Virtually every worker in our sample used Outlook regularly (6,996 out of 7,137) in the pre-period months and we suspect we are capturing a large share of work-related email activity.
In contrast, many sample workers show almost no Word activity during our pre-period months. Based on conversations with participating firms, we think this mostly reflects roles that do not involve document creation, though some low activity could reflect use of a different word proccessing tool. We do not expect that individual access to Copilot would shift many of these workers towards using Word, but if it did, that change from zero to positive activity would have a very different interpretation than the intensive margin behavioral changes of workers with recorded Word activity in the pre-period. We therefore restrict all Word-related analyses to workers with moderate use in the pre-period, indicated by above-median time spent in Word (0.25 hours/week) and above-median number of documents originated as the ‘primary editor’ (0.16 documents/week).
In the case of meetings, workers with near-zero pre-period use of Teams are concentrated in 10 firms, all of whom report primarily using non-Microsoft applications for video meetings. As with document creation, we are unable to establish any pre-period baseline meeting activity for these workers. In this case, we believe this largely reflects our inability to measure work meetings rather than true zeros. Any Copilot-induced changes in Teams activity is likely to reflect substitution to Teams rather than true changes in work meetings. We therefore drop workers at these 10 firms for all meeting outcome analyses.
@@ B.1 Outcome Metrics
We construct weekly metrics capturing each worker’s email, document editing, and meeting behaviors. Metrics are winsorized to the 99th percentile to address telemetry errors and privacy concerns.
@@
    For email, we aggregate the signals of various ‘reading actions’ individuals take into ‘Outlook sessions:’ blocks of time where they don’t go more than 15 minutes without reading an email. [^15] Read actions are the most commonly and frequently triggered action when people are interacting with Outlook and hence provide the best summary estimate of an Outlook session. Session time begins at the first eligible moment outside a prior session when a user opens an email in the reading pane. It ends either when the user exits the reading pane and doesn’t open an email for at least 15 minutes, with the exact end of the session being the earlier of the user’s exit or 15 minutes after the reading action.
@@
    • Outlook out-of-hours session time The subset of Outlook session time that occurred outside the individual’s “working day” as defined as the 8 hours in which they were most active during one fixed month of the pre-period immediately prior to the experiment.
@@
    • Unique email threads replied to The number of unique email threads/conversations among emails that are received that week and replied to within 7 days.
    • Time to reply from email delivery The average length of time it takes to reply to an email among emails received that week and replied to within 7 days.
@@
    • Total emails in threads replied to The total number of emails in the threads/conversations among emails that are received that week and replied to within 7 days.
@@
    • Teams meeting time The time spent in scheduled Teams meetings (excluding spontaneous calls) in a week, measured in hours.
    • Share of meetings attended late/early The share of scheduled Teams meetings joined either $\geq$5 minutes after the Teams meeting was started or left $\geq$5 minutes before the meeting was ended.
    • Recurring Teams meetings time The time spent in Teams meetings marked as recurring meetings, measured in hours.
@@
    Time use in Word is also measured by session time but is calculated based on intentional app actions (typing or changes to a document) rather than emails read individually or in sequence. Across metrics, a worker is a primary editor if they make the most edits on a document; a document is complete when 90% of the total edits to the document have been completed. [^16] Some documents show stray edits long after most active work is done, which may be accidental. This 90% cutoff aligns most often with other indicators of completion such as attaching to an email or printing to pdf.
    • Word total session time The time spent in Word sessions, identified as blocks of actions with $<15$ minute gaps, measured in hours.
@@
    • Avg. collaborative document time to complete The same as “avg. document time to complete” but only for documents with multiple collaborators.
@@ B.2 Coworkers
We also use the telemetry data to identify the closest coworkers of each person in the experiment, based on a weighted sum of their joint meetings, calls, emails, chats, and document collaborations. For each pair worker, we calculate each metric and divide by the firm-specific average. We add these normalized numbers together for an interaction score and call someone a “coworker” if the score is greater than 8.0. (Since there are five metrics, if a pair of workers have an entirely average level of interaction, their score would be 5.) If an individual has more than 15 coworkers, we take only the 15 with the highest interaction score. The average worker has 7 coworkers. We can then calculate what share of an individual’s coworkers have a Copilot license. Figure A2 shows the distribution of that share.
@@ C.1 Adoption
Figure A3 shows some additional Copilot usage patterns. Figure 3(a) shows the variation across firms in usage over time. Figure 3(b) shows a binscatter of the mean weekly Copilot actions taken in each app among users using some form of Copilot that week by the firm-level share of users that are active in an average week. Similar to Figure 1(b), we see that the usage intensity grows with firm usage for Teams, while it is flat or even declining for Outlook (with Word in between).
@@
Emails read -0.0003∗∗ -0.0002∗∗ -0.0002∗∗ -0.0003∗∗
($7.69\times 10^{-5}$) ($7.4\times 10^{-5}$) ($5.58\times 10^{-5}$) ($7.3\times 10^{-5}$)
Unique email threads replied to -0.0008 -0.0008 -0.0004 -0.0006
@@
Teams meeting total session time 0.0120∗∗ 0.0119∗∗ 0.0091∗∗ 0.0132∗∗
@@
Teams meetings attended 0.0072∗∗ 0.0073∗∗ 0.0035 0.0056
@@
Word total session time 0.0157 0.0161 0.0116 0.0202∗∗
@@
Others’ documents read/edited before completion 0.0422 0.0408 0.0325 0.0410
@@
Number of email recipients -0.0005 -0.0005 -0.0007 -0.0007
@@
Number of unique people met 1-1 (total) 0.0002 0.0002 0.0003 0.0002
@@
Share of coworkers with Copilot license 0.0302 0.0546∗∗ 0.0352
(0.0391) (0.0206) (0.0368)
Fixed-effects
Firm Yes Yes
Industry Yes Yes Yes
Post-period start month Yes
Fit statistics
Observations 3,422 3,422 3,684 3,422 3,684 3,684 3,422
R2 0.11840 0.11896 0.25608 0.29475 0.04462 0.11478 0.15394
Within R2 0.05014 0.11090
Significance: **$p<0.01$, *$p<0.05$
Note: This table shows OLS regressions results using individuals’ pre-experiment behaviors, share of coworkers with a Copilot license, and firm, industry, and post-period start month fixed effects to predict mean post-period Copilot weekly usage among treated workers. All variables are averages over the pre-period other than “number of unique people met 1-1,” which is a total over a fixed pre-period.
Table A3: Predicting Weekly Copilot Use
Emails read -0.0002∗∗ -0.0002∗∗ -0.0002∗∗ -0.0003∗∗
($7.3\times 10^{-5}$) ($6.99\times 10^{-5}$) ($5.64\times 10^{-5}$) ($6.96\times 10^{-5}$)
Unique email threads replied to -0.0008 -0.0008 -0.0004 -0.0006
@@
Teams meeting total session time 0.0132∗∗ 0.0131∗∗ 0.0092∗∗ 0.0140∗∗
@@
Teams meetings attended 0.0064∗∗ 0.0065∗∗ 0.0035 0.0050
@@
Word total session time 0.0158 0.0161 0.0115 0.0198∗∗
@@
Others’ documents read/edited before completion 0.0398 0.0389 0.0337 0.0393
@@
Number of email recipients -0.0005 -0.0005 -0.0007 -0.0006
@@
Number of unique people met 1-1 (total) 0.0003 0.0003 0.0003 0.0003
@@
Share of coworkers with Copilot license 0.0221 0.0535 0.0257
(0.0389) (0.0204) (0.0364)
@@
Firm Yes Yes
Industry Yes Yes Yes
Post-period start month Yes
@@ C.2 Outcomes
Note: These plots show the coefficients from regressions similar to Equation (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates. Vertical lines show 95% confidence intervals based on firm-clustered standard errors. Horizontal dashed lines give the average estimated effect from Table 2.
Table A4 shows the same analysis as Table 2 for some additional outcome variables. The decrease in emails read individually is unsurprising if Copilot is summarizing email threads (or all new emails) for people. Total emails replied to is another measure of reply activity where, while the difference is statistically significiant, it is practically very small ($<2\%$).
We also look at several different types of meetings. While we do see some increases for big meetings, the event studies in Figure A5 show that these increases only happened at the very start of the study period and seem likely due to the Copilot trainings that many companies conducted.
Table A4: Effects of Copilot on Additional Outcome Measures
@@
Total emails in email threads replied to 3.83 -0.038* -0.093* 7,095 320,811
@@
Total Teams meetings attended 9.93 0.20* 0.46 6,170 319,836
@@
Total recurring meetings 4.13 0.083* 0.19* 6,170 268,353
@@
Small short meetings 6.73 0.065 0.15 6,170 268,353
@@
Small long meetings 1.21 -0.001 -0.003 6,170 268,353
@@
Note: For each outcome variable, this table shows pre-period mean and cross-user standard deviation (Column 1), the intent-to-treat coefficients for the effect of Copilot (Column 2), and the LATE coefficients (Column 3). Cross-user standard deviations are estimates of the standard deviation of the random intercept in the model $Y_{it}=\mu_{i}+\epsilon_{it}$. Standard errors are clustered by firm, and asterisks are computed from sharpened $q$-values that account for the total number of hypotheses tested in the main text and appendices.
@@
Note: These plots show the coefficients from regressions similar to Equation (1) for the ITT (intent to treat) estimates and Equation (2) for the LATE estimates. Vertical lines show 95% confidence intervals based on firm-clustered standard errors. Horizontal dashed lines give the average estimated effect from Table A4.